OpenStudy (anonymous):

Consider the following. log4(x) (a) Rewrite the logarithm

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

Recall that \[log_4x =k \iff 4^k = x\] How are you wanting it written?

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

any way

OpenStudy (anonymous):

is x rite?

OpenStudy (anonymous):

What? I don't even understand what you're looking for here...

OpenStudy (anonymous):

Consider the following. log4(x) (a) Rewrite the logarithm

OpenStudy (anonymous):

i got anotehr problem similar to this one, look

OpenStudy (anonymous):

Rewrite it how?

OpenStudy (anonymous):

log8(x) (a) Rewrite the logarithm as a ratio of common logarithms

OpenStudy (anonymous):

Oh, sure.

OpenStudy (anonymous):

So what base do you want to use?

OpenStudy (anonymous):

Common bases are 10 or e.

OpenStudy (anonymous):

answer is log10(x/log10)(8)

OpenStudy (anonymous):

So you want to use 10 for the base?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Ok, so remember what I said.. \[log_4(x) = k \iff 4^k = x\] If \[4^k = x\]Then \[log_{10}(4^k) \ log_{10}(x)\]\[\implies k\cdot log_{10}(4) = log_{10}(x)\]\[\implies k = {log_{10}(x)\over log_{10}(4)}\] But since \(k = log_4(x)\) We have: \[log_4(x) = {log_{10}(x)\over log_{10}(4)}\]

OpenStudy (anonymous):

Log10(x)/log10(4) is your final answer?

OpenStudy (anonymous):

Log10(x)/log10(4) is your final answer?